# Beyond the EPICS: comprehensive Python IOC development with QueueIOC

### v2

@@ Abstract
Architectural deficiencies in EPICS lead to inefficiency in the development and application of EPICS input/output controllers (IOCs). An unintrusive solution is replacing EPICS IOCs with more maintainable and flexible Python IOCs, only reusing the Channel Access (CA) protocol of EPICS. After a digression about GUI development inspired by EPICS operator interfaces (OPIs), the structural similarity between standalone GUI backends, the Mamba backend, EPICS IOCs and other server programs is analysed. By combining the caproto library and event loops like those in these server programs, the QueueIOC framework for Python IOCs is created, which has the potential to systematically replace most EPICS IOCs currently used. Examples are first given for workalikes of StreamDevice and asyn; examples for seq-like applications include monochromators, motor anti-bumping and motor multiplexing. Also shown is software to use with the ~/iocBoot convention which addresses some issues with a similar solution based on procServ, along with a workalike of procServControl. A QueueIOC-based framework for detector integration, which overcomes areaDetector’s limitations in performance and architecture, is presented in an accompanying paper.
@@ 1 Introduction
The Experimental Physics and Industrial Control System (EPICS) is a basis for accelerator control and beamline control at HEPS, the High Energy Photon Source (Chu et al., 2018; Liu, Dong and Li, 2022). The rich selection of EPICS modules available has greatly facilitated device-control tasks at HEPS. However, during these tasks, it also became apparent that EPICS has some deficiencies inherent in its architecture (cf. Figure 4b), which lead to inefficiency in development and application.
@@
Our first complaints with EPICS are about its conception of “record links” (Kraimer et al., 2018), as well as the locking, scanning and processing mechanism associated with them. EPICS record links have different types and attributes: there are input links, output links and forward links; they can be subdivided into database links, Channel Access links and other subtypes; the “Process Passive” attribute can have the values NPP, PP, CA, CP and CPP. Events (hardware interrupts, user writing, periodic “scanning”) result in record “processing”, where the selection and ordering of records to be processed are determined by these links, according to the latter’s types and attributes. The rules for this selection and ordering are complex, and a major factor contributing to the complexity is the tight coupling between the types and attributes of record links. Ideally there should be separate rules about types, subtypes and attributes (especially “Process Passive” among the last) that add up to form the rules for individual combinations; the actual rules are quite far from able to be decomposed (the results would surely be easier to remember), and are instead quite like a lot of special cases piled together. The complexity of record links leads to considerable difficulty in learning: even among experienced EPICS application developers, few can precisely retell the rule for a less common combination of type and attribute. Fortunately, in practice only a handful of combinations are commonly used; on the other hand, even with the already complex set of rules, there are still requirements that cannot be naturally done, and a prime example for this is the motor module. In addition to the types and attributes of links, the selection and ordering of records to be processed are also affected by the records themselves, depending on them being “synchronous” or “asynchronous”; however, motor records are neither synchronous nor asynchronous.
@@
Apart from the link mechanism, EPICS records themselves are also quite inexpressive: EPICS “databases”, which are collections of records, can be nested; however, nested database files cannot be loaded at runtime, and instead must be expanded at compile time; .substitutions files can be loaded at runtime, but it cannot be nested either. This makes it very hard to abstract reusable functionalities as pluggable sub-databases, even for simple requirements like proportional-integral-derivative (PID) controllers. One idiomatic alternative is writing these functionalities as customised records, eg. EPID from the std module, which involves a considerable amount of boilerplate code and quite an extent of knowledge on EPICS internals. Another common alternative is writing “sequencers” based on the seq module, which is often error-prone because these “sequencers” can easily become comparable to C/C++ code riddled with goto statements; the monochromator code in the optics module are representative examples for this issue. The st.cmd language in EPICS does not have looping or conditional constructs, so we also cannot circumvent the inexpressiveness of records on the st.cmd layer. To work around the inexpressiveness of st.cmd itself, it is common practice to use external code generators, eg. iocbuilder (Abbott and Cobb, 2011), makeDb.py (from the ADGenICam module) and our own ones (cf. ADXspress3 in the next paragraph). We note that as can be seen from the examples given, code generators are in particular commonly used to deal with hardware that expose dynamic interfaces: variable numbers of detector channels, variable sets of camera features, etc. st.cmd also feels inconvenient to operators, as it is mainly intended as a developer-oriented language, and its inexpressiveness further impedes user-oriented abstraction.
@@
It is undoubted that none of the issues above is fatal to EPICS: developers can still produce usable, reliable EPICS modules, and users can still use these modules to actually fulfill their needs; as has been summarised above, the problems are about inexpressiveness and inconvenience. Meanwhile, during the construction of HEPS, we have found that improvements on expressiveness and convenience can be of great value in boosting efficiency; here we briefly give a few examples for this (Figure 1). The ihep-pkg packaging framework (Liu, Dong and Li, 2022, cf. also Section 4) now covers a large range of EPICS modules than that of the NSLS-II epicsdeb repository (Brookhaven National Laboratory, 2018), and provides support for CentOS 7, Rocky 8 plus partial support for Windows (using the MinGW environment), while still being easy to keep up to date. The ADXspress3 module (https://codeberg.org/CasperVector/ADXspress3) merges functionalities from both of its upstream versions (https://github.com/quantumdetectors/xspress3-epics and https://github.com/epics-modules/xspress3), and is much easier to maintain than both upstreams; meanwhile, to users it is also more convenient to build and use. The MambaPlanner mechanism (Li et al., 2023) provides a succinct, consistent command-line interface (CLI) for both step scans and fly scans, encapsulating hairy details about hardware configuration, correctness checks and data processing; while being convenient for debugging and advanced usage, it also facilitates development of graphical user interfaces (GUIs) oriented toward regular users. As is shown in Figure 1, they have each approached some kind of complexity lower-bound, in the sense that there is no obvious way to significantly simplify the code without inordinately complicating other code modules or sacrificing code clarity. Following the “succinctness is power” statement (Graham, 2002), by approaching the complexity lower-bound we can maximise the engineering efficiency. The examples above are just natural results of this rule; the efficiency boosts in them can be in 1–2 orders of magnitude, which also nicely coincide with the reduction of code needed to do the same tasks. Similarly, in other applications at HEPS, improvements in succinctness also prove to coincide with improvements in efficiency, even when the complexity lower-bounds have not yet been approached.
@@
From the perspective above, we can give a summary of the problems with EPICS: the essence of these problems is that its architecture makes the complexities of EPICS modules vastly higher than the lower bounds; the latter can often be estimated intuitively, just like in Figure 1. We understand that the design of EPICS is deeply affected by historical factors: the choice of record and links over a full-fledged language, as well as the inexpressiveness of st.cmd, may be largely seen as results of hardware limitations on Versa Module Europa (VME) and similar hardware platforms; record and links are also reminiscent of the programming paradigm common with programmable logical controllers (PLCs), PandABox (Christian et al., 2019) etc. Nevertheless, with the fast growth of EPICS usage on PCs, it is also time to rethink about these design decisions. Of course, an option for completely new facilities is to embrace ecosystems with less historical burdens, eg. Karabo (Göries et al., 2023); for facilities with more legacies, a more backward-compatible option is to reuse the Channel Access (CA) protocol of EPICS, but internally use more succinct tools. Given the common programming practices in last decade (especially those related to large scientific facilities), Python is an obvious candidate for this task; in particular, we have chosen the pure-Python caproto library (Allan, 2021), as it provides succinct interfaces for both server-side and client-side programming with the CA protocol. However, the officially expected usage of caproto depends on Python’s async/await programming paradigm, which leads to additional costs in learning and development; in this paper, we describe how we solve this problem, and produce a general-purpose workalike of EPICS based on Python.
@@ 2 The submit/notify pattern for GUI programming
As is hinted above, the main event loop in a GUI frontend should only contain logic and states essential to the graphical representation of the business logic – which in turn belongs to the backend; this improves the reliability and maintainability of GUIs, just like how privilege separation improves the security of qmail and software with similar designs (Bernstein, 2007). In the Mamba framework, the backend communicates with the frontends through a remote-procedure call (RPC) mechanism (Figure 2a) based on ZeroMQ. For tasks (eg. some in beamline control) that do not need the full Mamba infrastructure or even barebone Bluesky (Allan et al., 2019), we also developed some standalone GUIs with the submit/notify pattern; in these GUIs, there is still a clear separation between frontends and backends. Mamba frontends developed in the submit/notify style are, as of now, primarily used in our framework for automated attitude tuning of beamlines, presented in an accompanying paper (Li et al., 2024). The code for these frontends has been released in the open-source edition of Mamba (cf. Section 4); released in the same repository are two PyQt-based examples for standalone submit/notify GUIs. One of them is a workalike of the program in (Du et al., 2012), essentially an X-ray beam position monitor (XBPM) based on area detectors, with a simplified RPC mechanism between the backend and the frontend based on Python’s queue. The another is a simplified workalike of the ImageJ plugin provided by the ADViewers module, treating EPICS IOCs for area detectors as its backends, reusing the CA protocol for message passing with them.
@@ 3 The client-server model and EPICS IOCs
As has been discussed in Section 2, the CA protocol can be seen as a message-passing mechanism between servers (EPICS IOCs) and clients (EPICS OPIs etc); moreover, IOCs may be compared to GUI backends (whether standalone or in Mamba), and sometimes even be directly used as the backends. Therefore from analysing what is common in these servers, it is possible to find better ways to do what EPICS IOCs do; historically our standalone backends also learned from the Mamba backend, so here we begin with Mamba. Communication between the backend and frontends in Mamba may be categorised into requests/replies and notifications, the latter necessary due to the intrinsic weaknesses of state polling (Liu et al., 2022). In both standalone and Mamba backends there are main event loops: in Mamba backends, event handling (also including sending notifications) is separated into handlers registered to the core library, which encapsulates the main event loop; in standalone backends, there are explicit main loops that reply to requests and send notifications. We find the combination of an event loop, requests (with or without replies; an example for the latter is submissions in the submit/notify pattern, cf. Section 2) and notifications a common pattern in server programs. In addition to GUI backends, other examples are also abundant, eg. the alsamixer program to control audio volume under Linux (Figure 3): although not written with a frontend and a backend decoupled from each other, this program still features an event loop, which handles mouse/keyboard events and forwards status updates from the operating system. We are also aware of least one proprietary device-control product used at HEPS that explicitly uses requests/replies and notifications over TCP/IP network for the vendor’s application programming interface (API).
@@
In the light of these server programs, it is easy to find that caput, caget and camonitor, the basic operations in the CA protocol of EPICS, are also specialised combinations of requests/replies and notifications. Given these observations, we designed the basic architecture of our Python IOC framework, as in Figure 4(a). To accommodate the more mainstream programming style both inside and outside the Python ecosystem, the main event loop in the framework is currently based on a regular (non-async) function running in a regular Python thread. Correspondingly, there is a thin layer that isolates from the main loop the async/await-based code caproto expects (cf. Section 1); as the layer is based on variants of Python’s queue, we call our framework QueueIOC. For many IOCs in our framework the main loops are explicit, and are quite like the main loop in the standalone XBPM program in Section 2; however, for certain types of requirements exhibiting obvious regularity (eg. the QScanIOC-based IOCs in Section 4 and all IOCs in Section 5), we also encapsulate the boilerplate code in their main loops, so that the developer usually only needs to care about the essentials (cf. Figure 1, 5 & 8).
Figure 4: (a) The architecture of QueueIOC, in comparison with (b) the architecture of an EPICS IOC; the latter image is courtesy of Kraimer et al. (2018).
@@
Aside from QueueIOC and barebone caproto, we are also aware of similar attempts at using Python in the EPICS ecosystem, eg. PyDevice (Vodopivec, 2020) and pythonSoftIOC (Cobb, 2021). Instead of comparing them in detail, here we note the most crucial difference between QueueIOC and the rest is that it attempts to replace most EPICS IOCs in a systematic way, while striving to keep the amounts of code for the Python IOCs satisfactorily close to their intuitive complexity lower-bounds. For the latter goal, we explicitly note that QueueIOC focuses on fulfilling needs in succinct ways, which are not necessarily how EPICS fulfills the same needs. In the following sections, concrete examples will be given for various types of applications doable with QueueIOC; we believe these examples can adequately show the potential of QueueIOC. We also note that motor IOCs are perhaps a most obvious type of IOCs currently not covered by QueueIOC, but that we do not find this a fundamental weakness. This is because the implementation of the motor record contains a state machine with a few thousands lines of code, as well as a similar amount of ancillary code, which would take considerable human resource to port to Python even if we omitted some less useful features. A simple motor-like interface is indeed provided in QueueIOC, but this interface is, at least for now, intended only for “sequencers” (cf. Section 5) and not real motors.
@@
To fully understand the potential of QueueIOC, it is instructive to compare its architecture with the architecture of EPICS IOCs (Figure 4b). Database access, IOC database, record support and device support are either implicit or unnecessary in QueueIOC; device drivers and interrupt routines are implicit in the underlying control layer; monitors are implicit in the CA layer. By treating other IOCs (communicating through the CA protocol, cf. also Section 5) as the underlying control layer, “sequencers” are just a specialised kind of IOCs. With mechanisms like QScanIOC (cf. Section 4) and QSlowIOC (Zhang et al., 2024), what scanners do can also be done succinctly with QueueIOC. In summary, all architectural elements in EPICS IOCs have satisfactory counterparts in QueueIOC; so in the long term, we believe QueueIOC has the potential to be eventually capable of doing what is done with all EPICS IOCs currently used, except those IOCs that run under special resource constraints (most prominently VME IOCs). Furthermore, as has been noted above, caput, caget and camonitor are just specialised requests/replies and notifications; we think similar conclusions could be made about the PV Access (PVA) protocol added to EPICS in recent years. In conjunction with our observations about server programs, it is also a reasonable guess that similar conclusions could be made about the communication protocols in other device-control ecosystems, eg. TANGO and Karabo. If these guesses were sufficiently accurate, QueueIOC might become a starting point for some kind of unified device-control ecosystem, since QueueIOC intentionally encourages a programming style agnostic of the device-control ecosystem: the developer usually only needs to care about requests/replies and notifications – not caput, caget or camonitor.

### v3

@@ Abstract
Architectural deficiencies in EPICS lead to inefficiency in the development and application of EPICS input/output controllers (IOCs). An unintrusive solution is replacing EPICS IOCs with more maintainable and flexible Python IOCs, only reusing the Channel Access (CA) protocol of EPICS. After a digression about GUI development inspired by EPICS operator interfaces (OPIs), the structural similarity between standalone GUI backends, the Mamba backend, EPICS IOCs and other server-like programs is analysed. By combining the caproto library and event loops like those in these programs, the QueueIOC framework for Python IOCs is created, which has the potential to systematically replace most EPICS IOCs currently used. Examples are first given for workalikes of StreamDevice and asyn; examples for seq-like applications include monochromators, motor anti-bumping and motor multiplexing. Also shown is software to use with the ~/iocBoot convention which addresses some issues with a similar solution based on procServ, along with a workalike of procServControl. A QueueIOC-based framework for detector integration, which aims to overcome some architectural limitations of areaDetector while still offering decent performance, is presented in an accompanying paper.
@@ 1 Introduction
The Experimental Physics and Industrial Control System (EPICS) is a basis for accelerator control and beamline control at HEPS, the High Energy Photon Source (Chu et al., 2018; Liu, Dong and Li, 2022). The rich selection of EPICS modules available has greatly facilitated device-control tasks at HEPS. However, during these tasks, it also became apparent that EPICS has some deficiencies inherent in its architecture (cf. Figure 4b), which lead to inefficiency in development and application. Our first complaints with EPICS are about its conception of “record links” (Kraimer et al., 2018). EPICS record links have different types (input links, output links and forward links), subtypes (database links, Channel Access links etc) and attributes (eg. NPP, PP, CA, CP and CPP for the “Process Passive” attributes). Events (hardware interrupts, user writing, periodic “scanning”) result in record “processing”, where the selection and ordering of records to be processed are determined by these links according to a large and complex rule set. The rule set cannot be decomposed into separate rules about types, subtypes and attributes, which leads to considerable difficulty in learning. And even with the already complex rule set, there are still requirements that cannot be naturally done, and a prime example for this is the motor module. In addition to the types and attributes of links, the selection and ordering of records to be processed are also affected by the records themselves, depending on them being “synchronous” or “asynchronous”; however, motor records are neither synchronous nor asynchronous.
@@
Apart from the link mechanism, EPICS records themselves are also quite inexpressive: EPICS “databases”, which are collections of records, cannot be nested at runtime, which makes it very hard to abstract reusable functionalities as pluggable sub-databases, even for simple requirements like proportional-integral-derivative (PID) controllers. One idiomatic alternative is to implement customised records, eg. epid from the std module, which involves a considerable amount of boilerplate code and quite an extent of knowledge on EPICS internals. Another common alternative is writing “sequencers” based on the seq module, which is often error-prone (cf. Section 6). The st.cmd language in EPICS does not have looping or conditional constructs, so we also cannot circumvent the inexpressiveness of records on the st.cmd layer. A common practice to work around this is to use external code generators, eg. iocbuilder (Abbott and Cobb, 2011), makeDb.py (from the ADGenICam module) and our own ones (cf. ADXspress3 in the next paragraph). st.cmd also feels inconvenient to operators, further impeding user-oriented abstraction. It is undoubted that none of these issues is fatal to EPICS: developers can still produce usable, reliable EPICS modules, and users can still use these modules to actually fulfill their needs; as has been summarised above, the problems are about inexpressiveness and inconvenience. Meanwhile, during the construction of HEPS, we have found that improvements on expressiveness and convenience can be of great value in boosting efficiency; here we briefly give a few examples for this (Figure 1).
@@
The ihep-pkg packaging framework (Liu, Dong and Li, 2022, cf. also Section 4) now covers a larger range of EPICS modules than that of the NSLS-II epicsdeb repository (Brookhaven National Laboratory, 2018); it provides support for CentOS 7, Rocky 8 and hopefully compatible environments like RHEL 7/8, as well as partial support for Windows (using the MinGW environment); meanwhile it is still easy to keep up to date, even in comparison with recent works like installSynApps (Derbenev et al., 2023). The ADXspress3 module (https://codeberg.org/CasperVector/ADXspress3) merges functionalities from both of its upstream versions (https://github.com/quantumdetectors/xspress3-epics and https://github.com/epics-modules/xspress3), and is much easier to maintain than both upstreams; meanwhile, to users it is also more convenient to build and use. The MambaPlanner mechanism (Li et al., 2023) provides a succinct, consistent command-line interface (CLI) for both step scans and fly scans, encapsulating hairy details about hardware configuration, correctness checks and data processing; while being convenient for debugging and advanced usage, it also facilitates development of graphical user interfaces (GUIs) oriented toward regular users. As is shown in Figure 1, they have each approached some kind of complexity lower-bound, in the sense that there is no obvious way to significantly simplify the code without inordinately complicating other code modules or sacrificing code clarity. Following the “succinctness is power” statement (Graham, 2002), by approaching the complexity lower-bound we can maximise the engineering efficiency. The examples above are just natural results of this rule; the efficiency boosts in them can be in 1–2 orders of magnitude, which also nicely coincide with the reduction of code needed to do the same tasks. Similarly, in other applications at HEPS, improvements in succinctness also prove to coincide with improvements in efficiency, even when the complexity lower-bounds have not yet been approached.
@@
From the perspective above, we can give a summary of the problems with EPICS: the essence of these problems is that its architecture makes the complexities of EPICS modules vastly higher than the lower bounds; the latter can often be estimated intuitively, just like in Figure 1. We understand that the design of EPICS is deeply affected by historical factors: the choice of record and links over a full-fledged language, as well as the inexpressiveness of st.cmd, may be largely seen as results of hardware limitations on Versa Module Europa (VME) and similar hardware platforms; record and links are also reminiscent of the programming paradigm common with programmable logical controllers (PLCs), PandABox (Christian et al., 2019) etc. Nevertheless, with the fast growth of EPICS usage on PCs, it is also time to rethink about these design decisions. Of course, an option for completely new facilities is to embrace ecosystems with less historical burdens, eg. Karabo (Göries et al., 2023); for facilities with more legacies, a more backward-compatible option is to reuse the Channel Access (CA) protocol of EPICS, but internally use more succinct tools. Given the common programming practices in last decade (especially those related to large scientific facilities), Python is an obvious candidate for this task; in particular, we have chosen the pure-Python caproto library (Allan, 2021), as it provides succinct interfaces for both server-side and client-side programming with the CA protocol; alternatives like PCASPy (https://github.com/paulscherrerinstitute/pcaspy) may have issues like https://github.com/pyepics/pyepics/issues/176. However, the officially expected usage of caproto depends on Python’s async/await programming paradigm, which leads to additional costs in learning and development, although it is very helpful in highly concurrent application scenarios; in this paper, we describe how we solve this, and produce a general-purpose workalike of EPICS based on Python.
@@ 2 The submit/notify pattern for GUI programming
As is hinted above, the main event loop in a GUI frontend should only contain logic and states essential to the graphical representation of the business logic – which in turn belongs to the backend; this improves the reliability and maintainability of GUIs. In the Mamba framework, the backend communicates with the frontends through a remote-procedure call (RPC) mechanism (Figure 2a) based on ZeroMQ. For tasks (eg. some in beamline control) that do not need the full Mamba infrastructure or even barebone Bluesky (Allan et al., 2019), we also developed some standalone GUIs with the submit/notify pattern; in these GUIs, there is still a clear separation between frontends and backends. Mamba frontends developed in the submit/notify style are, as of now, primarily used in our framework for automated attitude tuning of beamlines, presented in an accompanying paper (Li et al., 2024). The code for these frontends has been released in the open-source edition of Mamba (cf. Section 4); released in the same repository are two PyQt-based examples for standalone submit/notify GUIs. One of them is a workalike of the program in (Du et al., 2012), essentially an X-ray beam-position monitor (XBPM) based on area detectors, with a simplified RPC mechanism between the backend and the frontend based on Python’s queue. The other is a simplified workalike of the ImageJ plugin provided by the ADViewers module, treating areaDetector IOCs as its backends, reusing the CA protocol for message passing with them.
@@ 3 The client-server model and EPICS IOCs
As has been discussed in Section 2, the CA protocol can be seen as a message-passing mechanism between servers (EPICS IOCs) and clients (EPICS OPIs etc); moreover, IOCs may be compared to GUI backends (whether standalone or in Mamba), and sometimes even be directly used as the backends. Therefore from analysing what is common in these server-like programs, it is possible to find better ways to do what EPICS IOCs do; historically our standalone backends also learned from the Mamba backend, so here we begin with Mamba. Communication between the backend and frontends in Mamba may be categorised into requests/replies and notifications, the latter necessary due to the intrinsic weaknesses of state polling (Liu et al., 2022). In both standalone and Mamba backends there are main event loops: in Mamba backends, event handling (also including sending notifications) is separated into handlers registered to the core library, which encapsulates the main event loop; in standalone backends, there are explicit main loops that reply to requests and send notifications. We find the combination of an event loop, requests (with or without replies; an example for the latter is submissions in the submit/notify pattern, cf. Section 2) and notifications a common pattern in server-like programs. In addition to GUI backends, other examples are also abundant, eg. the alsamixer program to control audio volume under Linux (Figure 3): although not written with a frontend and a backend decoupled from each other, this program still features an event loop, which handles mouse/keyboard events and forwards status updates from the operating system. We are also aware of least one proprietary device-control product used at HEPS that explicitly uses requests/replies and notifications over TCP/IP network for the vendor’s application programming interface (API).
@@
In the light of these server-like programs, it is easy to find that caput, caget and camonitor, the basic operations in the CA protocol of EPICS, are also specialised combinations of requests/replies and notifications. Given these observations, we designed the basic architecture of our Python IOC framework, as in Figure 4(a). To accommodate the more mainstream programming style both inside and outside the Python ecosystem, the main event loop in the framework is currently based on a regular (non-async) function running in a regular Python thread. Correspondingly, there is a thin layer that isolates from the main loop the async/await-based code caproto expects (cf. Section 1); as the layer is based on various message queues, we call our framework QueueIOC. For many IOCs in our framework the main loops are explicit, and are quite like the main loop in the standalone XBPM program in Section 2; however, for certain types of requirements exhibiting obvious regularity (eg. the QScanIOC-based IOCs in Section 4 and all IOCs in Section 5), we also encapsulate the boilerplate code in their main loops, so that the developer usually only needs to care about the essentials (cf. Figure 1, 5 & 8).
Figure 4: (a) The architecture of QueueIOC, in comparison with (b) the architecture of an EPICS IOC; the latter image is courtesy of Kraimer et al. (2018). We note that although (b) is for the stagnant 3.15.x branch of EPICS, its actively developed 7.x branch is not fundamentally different in terms of architecture (except for the introduction of the PVA protocol aside from CA), and is also affected by the issues summarised in this paper.
@@
Aside from QueueIOC, (barebone) caproto and PCASPy, we are also aware of other attempts at using Python in the EPICS ecosystem, eg. PyDevice (Vodopivec, 2020) and pythonSoftIOC (Cobb, 2021). Instead of comparing them in detail, here we note the most crucial difference between QueueIOC and the rest is that it attempts to replace most EPICS IOCs (in the narrow sense, structured like what makeBaseApp.pl creates) in a systematic way, while striving to keep the amounts of code for the Python IOCs satisfactorily close to their intuitive complexity lower-bounds. For the latter goal, we explicitly note that QueueIOC focuses on fulfilling needs in succinct ways, which are not necessarily how EPICS fulfills the same needs. In the following sections, concrete examples will be given for various types of applications doable with QueueIOC; we believe these examples can adequately show the potential of QueueIOC. Here we explicitly note that what we want is not to aggressively replace all existing EPICS IOCs, but instead to provide a smooth transition path to more efficient alternatives that are (protocol-wise) compatible. We also note that motor IOCs are perhaps a most obvious type of IOCs currently not covered by QueueIOC, but that we do not find this a fundamental weakness. This is because the implementation of the motor record contains a state machine with a few thousands lines of code, as well as a similar amount of ancillary code, which would take considerable human resource to port to Python even if we omitted some less useful features. A simple motor-like interface is indeed provided in QueueIOC, but this interface is, at least for now, intended only for “sequencers” (cf. Section 5) and not real motors.
@@
To fully understand the potential of QueueIOC, it is instructive to compare its architecture with the architecture of EPICS IOCs (Figure 4b). Database access, IOC database, record support and device support are either implicit or unnecessary in QueueIOC; device drivers and interrupt routines are implicit in the underlying control layer; monitors are implicit in the CA layer. By treating other IOCs (communicating through the CA protocol, cf. also Section 5) as the underlying control layer, “sequencers” are just a specialised kind of IOCs. With mechanisms like QScanIOC (cf. Section 4) and QSlowIOC (Zhang et al., 2024), what “scanners” do can also be done succinctly with QueueIOC. In summary, all architectural elements in EPICS IOCs have satisfactory counterparts in QueueIOC; so in the long term, we believe QueueIOC has the potential to be eventually capable of doing what is done with all EPICS IOCs currently used, except those IOCs that run under special resource constraints (most prominently VME IOCs). There may also be performance-related concerns about QueueIOC due to its Python-based nature, but we find it unnecessary to worry about: with the IOC qscan_rate available from the supplementary materials, it can be verified that QueueIOC can easily offer refresh (“scan”) rates up to at least 100 Hz and monitor rates up to at least 500 Hz, both of which are comparable to what EPICS IOCs are normally expected to do; with QDetectorIOC (Section 4), QueueIOC also offers decent detector readout performance comparable to areaDetector. Our outlook can also be extended beyond the CA protocol of EPICS: as has been noted above, caput, caget and camonitor are just specialised requests/replies and notifications; we think similar conclusions could be made about the PV Access (PVA) protocol added to EPICS in recent years. In conjunction with our observations about server-like programs, it is also a reasonable guess that similar conclusions could be made about the communication protocols in other device-control ecosystems, eg. TANGO and Karabo. If these guesses were sufficiently accurate, QueueIOC might become a starting point for some kind of unified device-control ecosystem, since QueueIOC intentionally encourages a programming style agnostic of the device-control ecosystem: the developer usually only needs to care about requests/replies and notifications – not caput, caget or camonitor.